# Arithmetic dynamics and Generalized Fermat’s conjecture

### v2

@@ 3. Generalized Fermat’s conjecture
Let $X:={\mathbb{P}}^{1}\times{\mathbb{P}}^{1}(=\operatorname{Proj}(K[X_{0},X_{1}])\times\operatorname{Proj}(K[Y_{0},Y_{1}]))$, and $f_{N}:X\to X$ $(N\geqslant 2$) is defined to be
$f_{N}((x_{0}:x_{1})\times(y_{0}:y_{1}))=(x_{0}^{N}:x_{1}^{N})\times(y_{0}^{N}:y_{1}^{N}).$
Let $p_{i}:X\to{\mathbb{P}}^{1}$ be the projection to the $i$-th factor, and $L=p_{1}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))\otimes p_{2}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))$. Note that $f_{N^{\prime}}\circ f_{N}=f_{N}\circ f_{N^{\prime}}=f_{NN^{\prime}}$ (multiplicative), and that $f_{N}:X\to X$ is compatible with respect to $L$, that is, $f_{N}^{*}(L)=L^{\otimes N}$. Let $Y$ be a subvariety given by
$aX_{0}Y_{0}+bX_{1}Y_{1}+cX_{0}Y_{1}+dX_{1}Y_{0}=(X_{0}\ Y_{0})A\begin{pmatrix}Y_{0}\\ Y_{1}\end{pmatrix}=0\quad(a,b,c,d\in K),$
where $A=\begin{pmatrix}a&c\\ d&b\end{pmatrix}$. Then $Y_{N}$ is given by
$aX_{0}^{N}Y_{0}^{N}+bX_{1}^{N}Y_{1}^{N}+cX_{0}^{N}Y_{1}^{N}+dX_{1}^{N}Y_{0}^{N}=(X^{N}_{0}\ Y^{N}_{0})A\begin{pmatrix}Y^{N}_{0}\\ Y^{N}_{1}\end{pmatrix}=0.$
We assume that $abcd\not=0$ and $\det A\not=0$. Then it is easy to see that $Y_{N}$ is smooth over $K$ for all NN, so its genus is $(N-1)^{2}$. Thus, (3) in Theorem 1.2 holds in this case.
Fix a positive integer $p\geqslant 2$. We set $g_{N}=f_{p^{N}}$ for $N\geqslant 1$. Then
$g_{N^{\prime}}\circ g_{N}=g_{N}\circ g_{N^{\prime}}=f_{p^{N}p^{N^{\prime}}}=f_{p^{N+N^{\prime}}}=g_{N+N^{\prime}},$
so $\{g_{N}\}_{N=1}^{\infty}$ is additive. Thus generalized Fermat’s conjecture holds for $\{g_{N}\}_{N=1}^{\infty}$.
Example 3.9.
Here we consider a family of Lattès maps. Let $E$ be an elliptic curve over $K$ and $\pi:E\to{\mathbb{P}}^{1}$ be the morphism such that $\deg(\pi)=2$ and $\pi(P)=\pi(-P)$, that is, ${\mathbb{P}}^{1}=E/\{\pm 1\}$. For an integer $N\geqslant 2$, let $\varphi_{N}:{\mathbb{P}}^{1}\to{\mathbb{P}}^{1}$ be the morphism such that the following diagram is commutative:
[…]
Then $\varphi_{N}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))={\mathcal{O}}_{{\mathbb{P}}^{1}}(N^{2})$ and $\varphi_{N}\circ\varphi_{N^{\prime}}=\varphi_{N^{\prime}}\circ\varphi_{N}=\varphi_{NN^{\prime}}$. Thus if we define $F_{N}:{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}\to{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ to be $f_{N}=\varphi_{N}\times\varphi_{N}$, then we have a system of endomorphisms compatible with $L=p_{1}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))\otimes p_{2}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))$, where $p_{i}:X\to{\mathbb{P}}^{1}$ is the projection to the $i$-th factor, so the same assertions hold as in Example 3.8.

### v3

@@ 4. Product case (multi-indexed version)
Let $f_{(A,A^{\prime})}$ $(A,A^{\prime}\geqslant 2$) be the endomorphism on ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}(=\operatorname{Proj}(K[X_{0},X_{1}])\times\operatorname{Proj}(K[Y_{0},Y_{1}]))$ defined to be
$f_{(A,A^{\prime})}((x_{0}:x_{1})\times(y_{0}:y_{1}))=(x_{0}^{A}:x_{1}^{A})\times(y_{0}^{A^{\prime}}:y_{1}^{A^{\prime}}).$
Let $p_{i}:{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}\to{\mathbb{P}}^{1}$ be the projection to the $i$-th factor, and $L=p_{1}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))\otimes p_{2}^{*}({\mathcal{O}}_{{\mathbb{P}}^{1}}(1))$. Note that $f_{(A,A^{\prime})}\circ f_{(B,B^{\prime})}=f_{(B,B^{\prime})}\circ f_{(A,A^{\prime})}=f_{(AB,A^{\prime}B^{\prime})}$ (multiplicative). Let $Y$ be a subvariety of ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ given by
$\alpha X_{0}Y_{0}+\beta X_{1}Y_{1}+\gamma X_{0}Y_{1}+\delta X_{1}Y_{0}=(X_{0}\ X_{1})\Pi\begin{pmatrix}Y_{0}\\ Y_{1}\end{pmatrix}=0\quad(\alpha,\beta,\gamma,\delta\in K),$
where $\Pi=\begin{pmatrix}\alpha&\gamma\\ \delta&\beta\end{pmatrix}$. Then $Y_{(A,A^{\prime})}$ is given by
$\alpha X_{0}^{A}Y_{0}^{A^{\prime}}+\beta X_{1}^{A}Y_{1}^{A^{\prime}}+\gamma X_{0}^{A}Y_{1}^{A^{\prime}}+\delta X_{1}^{A}Y_{0}^{A^{\prime}}=(X^{A}_{0}\ X^{A}_{1})\Pi\begin{pmatrix}Y^{A^{\prime}}_{0}\\ Y^{A^{\prime}}_{1}\end{pmatrix}=0.$
We assume that $\alpha\beta\gamma\delta\not=0$ and $\det\Pi\not=0$. Then it is easy to see that $Y_{(A,A^{\prime})}$ is smooth over $K$ for all $A,A^{\prime}\geqslant 2$, so its genus is $(A-1)(A^{\prime}-1)$. Thus, Theorem 4.5 holds in this case.